Not really sure on what I'm suppose to do. This is all I know.
PRACTICE: Use factoring to solve each equation for real values of the variable.
Please help me solve this equation: z^3/2 - z^1/2 = 0

Answer by Alan3354(69443)

z^3/2 - z^1/2 = 0
------------
Divide thru by z^(1/2)
z - 1 = 0
z = 1
By factoring:
z^(1/2)*(z - 1) = 0
z = 0, z = 1